Find the equation of the parabola with vertex (-2,0) and y intercept 4.
Can you please help me? Thanks so much in advance

Answer by lwsshak3(11628)

Find the equation of the parabola with vertex (-2,0) and y intercept 4.
**
Standard form of parabola: y=A(x-h)2+k, (h,k)=(x,y) coordinates of the vertex
For given problem:
y=A(x+2)^2+0
solve for A using coordinates of the y-intercept (0,4)
4=A(0+2)^2
4A=4
A=1
equation: y=(x+2)^2